# A Manifold Perspective on the Statistical Generalization of Graph Neural Networks

### v1

@@ Appendix C Local Lipschitz continuity of MNNs
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|\leq F^{L}C^{\prime}\|x-y\|,\quad\text{for all }x,y\in B_{r}({\mathcal{M}}),$ (60)
@@
$\leq\left|\sum_{q=1}^{F}{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(x)-\sum_{q=1}^{F}{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(y)\right|\leq F\left|f_{L-1}^{q}(x)-f_{L-1}^{q}(y)\right|.$ (62)
@@
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|\leq F^{L}|f(x)-f(y)|.$ (63)
With $f$ as a $\lambda$-bandlimited manifold signal, we suppose it possesses $M$ spectral components. The gradient can be bounded according to [52] as
$\|\nabla f\|_{\infty}\leq C\sum_{\lambda_{i}\leq\lambda}\lambda_{i}\|\bm{\phi}_{i}\|_{\infty}$ (64)
From Theorem 4.5 in [53], $f$ is locally Lipschitz continuous as
$|f(x)-f(y)|\leq C^{\prime}\|x-y\|,\quad\text{with }x,y\in B_{r}({\mathcal{M}}),$ (65)
@@
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|\leq F^{L}C^{\prime}\|x-y\|,\quad\text{with }x,y\in B_{r}({\mathcal{M}}),$ (66)
@@ Appendix D Proof of Theorem 1
To analyze the difference between the empirical risk and statistical risk, we introduce an intermediate term which is the induced version of the sampled MNN output. We define ${\mathbf{I}}_{N}$ as the inducing operator based on the Voronoi decomposition $\{V_{i}\}_{i=1}^{N}$ defined in (21). This intermediate term is written explicitly as
@@
Suppose ${\mathbf{H}}^{*}\in\arg\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathcal{M}}({\mathbf{H}})$, we have
$GA=\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathbf{G}}({\mathbf{H}})-\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathcal{M}}({\mathbf{H}})\leq R_{\mathbf{G}}({\mathbf{H}}^{*})-R_{\mathcal{M}}({\mathbf{H}}^{*})$ (68)
The difference between $R_{\mathbf{G}}({\mathbf{H}}^{*})$ and $R_{\mathcal{M}}({\mathbf{H}}^{*})$ can be decomposed as
$R_{\mathbf{G}}({\mathbf{H}}^{*})-R_{\mathcal{M}}({\mathbf{H}}^{*})$
$=\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)$ (69)
$=\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)$
$\qquad\qquad+\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)$ (70)
$\leq\left|\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$
$\qquad\qquad+\left|\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (71)
@@
$\left|\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (72)
$=\frac{1}{N}\left|\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\sum_{i=1}^{N}\ell(\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x_{i}),g(x_{i}))\right|$ (73)
$\leq\frac{1}{N}\sum_{i=1}^{N}\left|\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\ell(\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x_{i}),g(x_{i}))\right|$ (74)
$\leq\frac{1}{N}\sum_{i=1}^{N}\Big{|}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}})]_{i}-\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x_{i})\Big{|}$ (75)
$\leq\frac{1}{N}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},{\mathbf{I}}_{N}{\mathbf{x}})\|_{1}$ (76)
@@
$\left|\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (78)
$\leq\left|\sum_{i=1}^{N}\int_{V_{i}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\mathbbm{1}_{x\in V_{i}}\text{d}\mu(x)-\sum_{i=1}^{N}\int_{V_{i}}\ell\left({\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (79)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\ell\left(\overline{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)-\ell\left({\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\right|\mathbbm{1}_{x\in V_{i}}\text{d}\mu(x)$ (80)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\overline{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x)-{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x)\right|\mathbbm{1}_{x\in V_{i}}\text{d}\mu(x)$ (81)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x_{i})-\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x)\right|\mathbbm{1}_{x\in V_{i}}\text{d}\mu(x)$ (82)
From (78) to (79), it relies on the decomposition of the MNN output over $\{V_{i}\}_{i=1}^{N}$. From (80) to (82), we use the Lipschitz continuity of loss function. Proposition 3 indicates that the MNN outputs are Lipschitz continuous within a certain range, which leads to
$\sum_{i=1}^{N}\int_{V_{i}}\left|\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x_{i})-\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f)(x)\right|\mathbbm{1}_{x\in V_{i}}\text{d}\mu(x)\leq F^{L}C_{3}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}.$ (83)
Combining (77) and (83), we can conclude the proof. ∎
@@ Appendix E Proof of Theorem 2
Suppose ${\mathbf{H}}^{*}\in\arg\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathcal{M}}({\mathbf{H}})$, we have
$GA=\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathbf{G}}({\mathbf{H}})-\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathcal{M}}({\mathbf{H}})\leq R_{\mathbf{G}}({\mathbf{H}}^{*})-R_{\mathcal{M}}({\mathbf{H}}^{*})$ (84)
@@
$R_{\mathbf{G}}({\mathbf{H}}^{*})-R_{\mathcal{M}}({\mathbf{H}}^{*})$
$=\sum_{k=1}^{K}\left(\ell\left(\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i},y_{k}\right)-\ell\left(\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x),y_{k}\right)\right)$ (85)
Based on the property of absolute value inequality, we have
$\ell\left(\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i},y_{k}\right)-\ell\left(\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x),y_{k}\right)$
$\leq\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (86)
We insert an intermediate term $\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})$ as the value evaluated on the sampled point $x_{i}$, which leads to
$\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (87)
$\leq\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})\right|+$
$\qquad\qquad\qquad\qquad\qquad\qquad\left|\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (88)
@@
$\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})\right|$ (89)
$\leq\frac{1}{N}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}}_{k})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f_{k})\|_{1}$ (90)
$\leq\frac{1}{\sqrt{N}}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}}_{k})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},f_{k})\|_{2}$ (91)
@@
$\left|\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (93)
$=\left|\sum_{i=1}^{N}\int_{V_{i}}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})\mathbbm{1}_{x\in V_{i}}\text{d}\mu_{k}(x)-\sum_{i=1}^{N}\int_{V_{i}}\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})(x)\text{d}\mu_{k}(x)\right|$ (94)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})-\bm{\Phi}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho,k},f_{k})(x)\right|\mathbbm{1}_{x\in V_{i}}\text{d}\mu_{k}(x)$ (95)

### v2

@@ Appendix C Local Lipschitz continuity of MNNs
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|\leq F^{L}C^{\prime}\text{dist}(x-y),\quad\text{for all }x,y\in B_{r}({\mathcal{M}}),$ (70)
@@
$\leq\left|\sum_{q=1}^{F}{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(x)-\sum_{q=1}^{F}{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(y)\right|\leq F\max_{q=1,\cdots,F}\left|{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(x)-{\mathbf{h}}_{L}^{q}({\mathcal{L}}_{\rho})f_{L-1}^{q}(y)\right|.$ (72)
@@
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|\leq F^{L}|{\mathbf{h}}_{L}({\mathcal{L}}_{\rho})\cdots{\mathbf{h}}_{1}({\mathcal{L}}_{\rho})f(x)-{\mathbf{h}}_{L}({\mathcal{L}}_{\rho})\cdots{\mathbf{h}}_{1}({\mathcal{L}}_{\rho})f(y)|.$ (73)
With $f$ as a $\lambda$-bandlimited manifold signal, we suppose $g=bbh_{L}({\mathcal{L}}_{\rho})\cdots{\mathbf{h}}_{1}({\mathcal{L}}_{\rho})f$. As $\langle f,\bm{\phi}_{i}\rangle=0$ for all $i>M$, $g$ is also bandlimited and possesses $M$ spectral components. The gradient can be bounded according to [51] combined with the non-amplifying property of the filter function as
$\|\nabla g\|_{\infty}\leq C\sum_{\lambda_{i}\leq\lambda}\left|\hat{h}(\lambda_{i})\right|^{L}\lambda_{i}^{\frac{d+1}{2}}\|f\|_{\mathcal{M}}\leq C\sum_{\lambda_{i}\leq\lambda}\lambda_{i}^{\frac{d+1}{2}}\|f\|_{\mathcal{M}}$ (74)
From Theorem 4.5 in [52], $g$ is locally Lipschitz continuous as
$|g(x)-g(y)|\leq C^{\prime}\text{dist}(x-y),\quad\text{with }x,y\in B_{r}({\mathcal{M}}),$ (75)
@@
$|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(y)|\leq F^{L}C^{\prime}\text{dist}(x-y),\quad\text{with}x,y\in B_{r}({\mathcal{M}}),$ (76)
@@ Appendix D Proof of Theorem 1
To analyze the difference between the empirical risk and statistical risk, we introduce an intermediate term which is the induced version of the sampled MNN output. We define ${\mathbf{I}}_{N}$ as the inducing operator based on the Voronoi decomposition $\{V_{i}\}_{i=1}^{N}$ defined in Section A. This intermediate term is written explicitly as
@@
Suppose ${\mathbf{H}}\in\arg\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathcal{M}}({\mathbf{H}})$, we have
$GA=\sup_{{\mathbf{H}}\in{\mathcal{H}}}|R_{\mathbf{G}}({\mathbf{H}})-R_{\mathcal{M}}({\mathbf{H}})|$ (78)
The difference between $R_{\mathbf{G}}({\mathbf{H}})$ and $R_{\mathcal{M}}({\mathbf{H}})$ can be decomposed as
$\left|R_{\mathbf{G}}({\mathbf{H}})-R_{\mathcal{M}}({\mathbf{H}})\right|$
$=\left|\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (79)
$=\Bigg{|}\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)$
$\qquad\qquad+\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\Bigg{|}$ (80)
$\leq\left|\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$
$\qquad\qquad+\left|\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (81)
@@
$\left|\frac{1}{N}\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (82)
$=\frac{1}{N}\left|\sum_{i=1}^{N}\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\sum_{i=1}^{N}\ell(\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i}),g(x_{i}))\right|$ (83)
$\leq\frac{1}{N}\sum_{i=1}^{N}\left|\ell([\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i},[{\mathbf{y}}]_{i})-\ell(\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i}),g(x_{i}))\right|$ (84)
$\leq\frac{1}{N}\sum_{i=1}^{N}\Big{|}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})]_{i}-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i})\Big{|}$ (85)
$\leq\frac{1}{N}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},{\mathbf{I}}_{N}{\mathbf{x}})\|_{1}$ (86)
$\leq\frac{1}{\sqrt{N}}LF^{L-1}\left(C_{1}\epsilon+C_{2}\sqrt{\frac{\log(1/\delta)}{N}}\right)$ (87)
@@
$\left|\int_{\mathcal{M}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\int_{\mathcal{M}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (88)
$\leq\left|\sum_{i=1}^{N}\int_{V_{i}}\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)-\sum_{i=1}^{N}\int_{V_{i}}\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\text{d}\mu(x)\right|$ (89)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\ell\left(\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)-\ell\left({\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x),{g}(x)\right)\right|\text{d}\mu(x)$ (90)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)-{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)\right|\text{d}\mu(x)$ (91)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i})-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)\right|\text{d}\mu(x)$ (92)
From (88) to (89), it relies on the decomposition of the MNN output over $\{V_{i}\}_{i=1}^{N}$. From (90) to (91), we use the Lipschitz continuity of loss function. From (91) to (92), we use the definition of $\overline{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f)$. Proposition 3 indicates that the MNN outputs are Lipschitz continuous within a certain range, which leads to
$\sum_{i=1}^{N}\int_{V_{i}}\left|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x_{i})-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)(x)\right|\text{d}\mu(x)$
$\leq\sum_{i=1}^{N}\int_{V_{i}}F^{L}C_{3}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}\text{d}\mu(x)$ (93)
$=F^{L}C_{3}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}\sum_{i=1}^{N}\int_{V_{i}}\text{d}\mu(x)$ (94)
$\leq F^{L}C_{3}\left(\frac{\log N}{N}\right)^{\frac{1}{d}}.$ (95)
Combining (87) and (D), we can conclude the proof. ∎
@@ Appendix E Proof of Theorem 2
$\left|R_{\mathbf{G}}({\mathbf{H}})-R_{\mathcal{M}}({\mathbf{H}})\right|$
$\leq\sum_{k=1}^{K}\left|\ell\left(\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i},y_{k}\right)-\ell\left(\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x),y_{k}\right)\right|$ (96)
Based on the property of absolute value inequality and the Lipschitz continuity assumption of loss function (Assumption 2), we have
$\left|\ell\left(\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i},y_{k}\right)-\ell\left(\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x),y_{k}\right)\right|$
$\leq\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (97)
We insert an intermediate term $\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})$ as the value evaluated on the sampled point $x_{i}$, which leads to
$\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (98)
$\leq\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})\right|+$
$\qquad\qquad\qquad\qquad\qquad\qquad\left|\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (99)
@@
$\left|\frac{1}{N}\sum_{i=1}^{N}[\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N,k},{\mathbf{x}}_{k})]_{i}-\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})\right|$ (100)
$\leq\frac{1}{N}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}}_{k})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f_{k})\|_{1}$ (101)
$\leq\frac{1}{\sqrt{N}}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}}_{k})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},f_{k})\|_{2}$ (102)
@@
$\left|\frac{1}{N}\sum_{i=1}^{N}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})-\int_{{\mathcal{M}}_{k}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})\text{d}\mu_{k}(x)\right|$ (104)
$=\left|\sum_{i=1}^{N}\int_{V_{i}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})\text{d}\mu_{k}(x)-\sum_{i=1}^{N}\int_{V_{i}}\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x)\text{d}\mu_{k}(x)\right|$ (105)
$\leq\sum_{i=1}^{N}\int_{V_{i}}\left|\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x_{i})-\bm{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho,k},f_{k})(x)\right|\text{d}\mu_{k}(x)$ (106)